OpenStudy (anonymous):

The half-life of a certain radioactive material is 37 days. An initial amount of the material has a mass of 477 kg. Write an exponential function that models the decay of this material. Find how much radioactive material remains after 6 days. Round your answer to the nearest thousandth. It's multiple choice but in not sure how to type in the answers..

OpenStudy (anonymous):

a) b) c) d)

OpenStudy (anonymous):

i have to have them to help you

OpenStudy (anonymous):

A. \[y= 477 (1/2)^37x ; 0 kg \] Its ^37x Thats all sopposed to be small but i dont know how to do that. B. \[y=477(1/2)^1/37x ; 426.288kg\] Again its ^1/37 x all small C. \[y=2 (1/477)^1/37x ; 0.736 kg\] Same again. ^1/37x D. \[y=1/2 (1/477)^1/37x ;0.184kg\] ^1/37x All supposed to be small. Like part of the exponent.

OpenStudy (anonymous):

i would use \[477\times \left(\frac{1}{2}\right)^{\frac{t}{37}}\]

OpenStudy (anonymous):

to answer the second part, replace \(t\) by \(6\) and use a calculator

OpenStudy (anonymous):

Thanks! (:

OpenStudy (anonymous):

yw